# Training Hybrid Deep Quantum Neural Networks for Efficient Reinforcement Learning

### v5

@@ B.6 Backward pass of hDQNN–TD3 (update phase)
On the same minibatch we refine the surrogate by minimising the BCE loss of Eq. (5):
@@
Appendix C Additional graphs and tables
C.1 Extended Ablation
@@
4 00040004\,0004 000 FC 0 0 $423.0\pm 170.8$ $603.6$ $31.5$ 214 686214686214\,686214 686 0,1,2,3,4
4 00040004\,0004 000 RBG 0 0 $333.6\pm 196.5$ $545.3$ $35.7$ 198 449198449198\,449198 449 0,1,2,3,4
4 00040004\,0004 000 PQC $100$ 5 1 575.9±1 923.4plus-or-minus1575.91923.41\,575.9$\pm$ 1\,923.41 575.9 ± 1 923.4 4 882.74882.74\,882.74 882.7 $515.2$ 540 138540138540\,138540 138 0,1,2,3,4
4 00040004\,0004 000 PQC 1 00010001\,0001 000 5 $476.5\pm 149.4$ $650.8$ $225.3$ 280 476280476280\,476280 476 0,1,2,3,4
4 00040004\,0004 000 PQC $010$ $667.3\pm 442.1$ 1 453.31453.31\,453.31 453.3 $315.2$ 335 815335815335\,815335 815 0,1,2,3,4
4 00040004\,0004 000 PQC 1 $001 0$ $556.0\pm 81.7$ $689.4$ $265.5$ 308 709308709308\,709308 709 0,1,2,3,4
6 20062006\,2006 200 FC 0 0 $543.4\pm 243.3$ $909.4$ $72.2$ 423 010423010423\,010423 010 0,1,2,3,4
6 20062006\,2006 200 RBG 0 0 $417.0\pm 131.4$ $526.2$ $75.5$ 366 904366904366\,904366 904 0,1,2,3,4
6 20062006\,2006 200 PQC $100$ 5 2 014.7±2 008.6plus-or-minus2014.72008.62\,014.7$\pm$ 2\,008.62 014.7 ± 2 008.6 4 342.44342.44\,342.44 342.4 1 596.61596.61\,596.61 596.6 1 387 70513877051\,387\,7051 387 705 0,1,2,3,4
6 20062006\,2006 200 PQC 1 00010001\,0001 000 5 $544.8\pm 395.1$ 1 102.71102.71\,102.71 102.7 $467.0$ 517 611517611517\,611517 611 0,1,2,3,4
6 20062006\,2006 200 PQC $010$ 1 492.3±2 191.2plus-or-minus1492.32191.21\,492.3$\pm$ 2\,191.21 492.3 ± 2 191.2 5 411.25411.25\,411.25 411.2 $988.9$ 959 152959152959\,152959 152 0,1,2,3,4
6 20062006\,2006 200 PQC 1 $001 0$ $657.0\pm 145.0$ $841.6$ $552.6$ 610 225610225610\,225610 225 0,1,2,3,4
7 00070007\,0007 000 FC 0 0 $693.3\pm 453.1$ 1 473.01473.01\,473.01 473.0 $98.5$ 515 399515399515\,399515 399 0,1,2,3,4
7 00070007\,0007 000 RBG 0 0 $466.7\pm 106.0$ $547.2$ $92.8$ 438 296438296438\,296438 296 0,1,2,3,4
7 00070007\,0007 000 PQC $100$ 5 2 139.6±2 180.2plus-or-minus2139.62180.22\,139.6$\pm$ 2\,180.22 139.6 ± 2 180.2 4 962.04962.04\,962.04 962.0 1 935.91935.91\,935.91 935.9 1 756 26617562661\,756\,2661 756 266 0,1,2,3,4
7 00070007\,0007 000 PQC 1 00010001\,0001 000 5 $733.1\pm 740.4$ 1 974.31974.31\,974.31 974.3 $612.5$ 642 867642867642\,867642 867 0,1,2,3,4
7 00070007\,0007 000 PQC $010$ $532.7\pm 105.6$ $690.6$ $678.9$ 666 056666056666\,056666 056 0,1,2,4
7 00070007\,0007 000 PQC 1 $001 0$ $710.1\pm 238.6$ 1 040.61040.61\,040.61 040.6 $675.5$ 744 615744615744\,615744 615 0,1,2,3,4
7 20072007\,2007 200 FC 0 0 $617.5\pm 341.1$ 1 207.81207.81\,207.81 207.8 $108.8$ 544 790544790544\,790544 790 0,1,2,3,4
7 20072007\,2007 200 RBG 0 0 $414.3\pm 103.0$ $545.0$ $99.6$ 456 050456050456\,050456 050 0,1,2,3,4
7 20072007\,2007 200 PQC $100$ 5 1 605.0±2 202.2plus-or-minus1605.02202.21\,605.0$\pm$ 2\,202.21 605.0 ± 2 202.2 4 906.64906.64\,906.64 906.6 1 791.91791.91\,791.91 791.9 1 527 03715270371\,527\,0371 527 037 0,1,2,4
7 20072007\,2007 200 PQC 1 00010001\,0001 000 5 1 078.5±1 517.3plus-or-minus1078.51517.31\,078.5$\pm$ 1\,517.31 078.5 ± 1 517.3 3 759.23759.23\,759.23 759.2 $673.0$ 692 138692138692\,138692 138 0,1,2,3,4
7 20072007\,2007 200 PQC $010$ $517.8\pm 95.0$ $659.2$ $712.2$ 694 104694104694\,104694 104 0,1,2,4
7 20072007\,2007 200 PQC 1 $001 0$ $763.5\pm 335.7$ 1 277.41277.41\,277.41 277.4 $711.9$ 782 646782646782\,646782 646 0,1,2,3,4
7 70077007\,7007 700 FC 0 0 1 126.1±1 467.2plus-or-minus1126.11467.21\,126.1$\pm$ 1\,467.21 126.1 ± 1 467.2 3 745.53745.53\,745.53 745.5 $135.4$ 621 311621311621\,311621 311 0,1,2,3,4
7 70077007\,7007 700 RBG 0 0 $432.4\pm 97.4$ $543.2$ $110.8$ 500 487500487500\,487500 487 0,1,2,3,4
7 70077007\,7007 700 PQC $100$ 5 1 402.8±1 757.3plus-or-minus1402.81757.31\,402.8$\pm$ 1\,757.31 402.8 ± 1 757.3 4 031.14031.14\,031.14 031.1 1 943.21943.21\,943.21 943.2 1 688 88816888881\,688\,8881 688 888 0,1,2,4
7 70077007\,7007 700 PQC 1 00010001\,0001 000 5 1 121.4±1 558.9plus-or-minus1121.41558.91\,121.4$\pm$ 1\,558.91 121.4 ± 1 558.9 3 868.63868.63\,868.63 868.6 $839.3$ 824 076824076824\,076824 076 0,1,2,3,4
7 70077007\,7007 700 PQC $010$ $506.6\pm 97.0$ $624.5$ $793.6$ 765 043765043765\,043765 043 0,1,2,4
7 70077007\,7007 700 PQC 1 $001 0$ $891.8\pm 547.1$ 1 852.91852.91\,852.91 852.9 $805.0$ 882 668882668882\,668882 668 0,1,2,3,4
@@
4 00040004\,0004 000 FC 0 0 $382.7\pm 167.5$ $603.6$ $27.4$ 196 401196401196\,401196 401 0,1,2,4
4 00040004\,0004 000 RBG 0 0 $356.5\pm 219.1$ $545.3$ $39.8$ 210 584210584210\,584210 584 0,1,2,4
4 00040004\,0004 000 PQC $100$ 5 1 555.4±1 923.4plus-or-minus1555.41923.41\,555.4$\pm$ 1\,923.41 555.4 ± 1 923.4 4 882.74882.74\,882.74 882.7 $564.8$ 574 890574890574\,890574 890 0,1,2,4
4 00040004\,0004 000 PQC 1 00010001\,0001 000 5 $521.3\pm 128.0$ $650.8$ $237.5$ 302 738302738302\,738302 738 0,1,2,4
4 00040004\,0004 000 PQC $010$ $470.8\pm 57.1$ $534.5$ $291.2$ 302 948302948302\,948302 948 0,1,2,4
4 00040004\,0004 000 PQC 1 $001 0$ $568.0\pm 89.2$ $689.4$ $258.4$ 303 057303057303\,057303 057 0,1,2,4
6 20062006\,2006 200 RBG 0 0 $389.7\pm 134.3$ $521.2$ $80.4$ 372 226372226372\,226372 226 0,1,2,4
6 20062006\,2006 200 FC 0 0 $514.3\pm 270.8$ $909.4$ $61.1$ 384 622384622384\,622384 622 0,1,2,4
6 20062006\,2006 200 PQC $100$ 5 1 432.7±1 766.9plus-or-minus1432.71766.91\,432.7$\pm$ 1\,766.91 432.7 ± 1 766.9 4 075.34075.34\,075.34 075.3 1 489.31489.31\,489.31 489.3 1 196 50311965031\,196\,5031 196 503 0,1,2,4
6 20062006\,2006 200 PQC 1 00010001\,0001 000 5 $602.9\pm 430.8$ 1 102.71102.71\,102.71 102.7 $498.4$ 567 036567036567\,036567 036 0,1,2,4
6 20062006\,2006 200 PQC $010$ $512.5\pm 52.6$ $590.5$ $551.2$ 557 415557415557\,415557 415 0,1,2,4
6 20062006\,2006 200 PQC 1 $001 0$ $652.4\pm 167.0$ $841.6$ $544.0$ 593 357593357593\,357593 357 0,1,2,4
7 00070007\,0007 000 FC 0 0 $691.6\pm 523.2$ 1 473.01473.01\,473.01 473.0 $81.9$ 467 999467999467\,999467 999 0,1,2,4
7 00070007\,0007 000 RBG 0 0 $446.5\pm 110.8$ $543.4$ $98.2$ 440 609440609440\,609440 609 0,1,2,4
7 00070007\,0007 000 PQC $100$ 5 1 666.9±2 201.8plus-or-minus1666.92201.81\,666.9$\pm$ 2\,201.81 666.9 ± 2 201.8 4 962.04962.04\,962.04 962.0 1 730.91730.91\,730.91 730.9 1 460 22814602281\,460\,2281 460 228 0,1,2,4
7 00070007\,0007 000 PQC 1 00010001\,0001 000 5 $837.8\pm 811.1$ 1 974.31974.31\,974.31 974.3 $664.2$ 711 615711615711\,615711 615 0,1,2,4
7 00070007\,0007 000 PQC $010$ $532.7\pm 105.6$ $690.6$ $678.9$ 666 056666056666\,056666 056 0,1,2,4
7 00070007\,0007 000 PQC 1 $001 0$ $698.3\pm 273.8$ 1 040.61040.61\,040.61 040.6 $665.6$ 724 890724890724\,890724 890 0,1,2,4
7 20072007\,2007 200 FC 0 0 $620.8\pm 393.8$ 1 207.81207.81\,207.81 207.8 $89.8$ 494 098494098494\,098494 098 0,1,2,4
7 20072007\,2007 200 RBG 0 0 $426.8\pm 114.5$ $545.0$ $105.9$ 458 399458399458\,399458 399 0,1,2,4
7 20072007\,2007 200 PQC $100$ 5 1 605.0±2 202.2plus-or-minus1605.02202.21\,605.0$\pm$ 2\,202.21 605.0 ± 2 202.2 4 906.64906.64\,906.64 906.6 1 791.91791.91\,791.91 791.9 1 527 03715270371\,527\,0371 527 037 0,1,2,4
7 20072007\,2007 200 PQC 1 00010001\,0001 000 5 1 269.5±1 681.1plus-or-minus1269.51681.11\,269.5$\pm$ 1\,681.11 269.5 ± 1 681.1 3 759.23759.23\,759.23 759.2 $735.6$ 770 278770278770\,278770 278 0,1,2,4
7 20072007\,2007 200 PQC $010$ $517.8\pm 95.0$ $659.2$ $712.2$ 694 104694104694\,104694 104 0,1,2,4
7 20072007\,2007 200 PQC 1 $001 0$ $768.9\pm 387.4$ 1 277.41277.41\,277.41 277.4 $700.9$ 761 498761498761\,498761 498 0,1,2,4
7 70077007\,7007 700 FC 0 0 1 253.1±1 662.1plus-or-minus1253.11662.11\,253.1$\pm$ 1\,662.11 253.1 ± 1 662.1 3 745.53745.53\,745.53 745.5 $115.1$ 573 588573588573\,588573 588 0,1,2,4
7 70077007\,7007 700 RBG 0 0 $431.2\pm 112.4$ $543.2$ $118.3$ 504 233504233504\,233504 233 0,1,2,4
7 70077007\,7007 700 PQC $100$ 5 1 402.8±1 757.3plus-or-minus1402.81757.31\,402.8$\pm$ 1\,757.31 402.8 ± 1 757.3 4 031.14031.14\,031.14 031.1 1 943.21943.21\,943.21 943.2 1 688 88816888881\,688\,8881 688 888 0,1,2,4
7 70077007\,7007 700 PQC 1 00010001\,0001 000 5 1 326.7±1 720.3plus-or-minus1326.71720.31\,326.7$\pm$ 1\,720.31 326.7 ± 1 720.3 3 868.63868.63\,868.63 868.6 $933.4$ 927 753927753927\,753927 753 0,1,2,4
7 70077007\,7007 700 PQC $010$ $506.6\pm 97.0$ $624.5$ $793.6$ 765 043765043765\,043765 043 0,1,2,4
7 70077007\,7007 700 PQC 1 $001 0$ $945.3\pm 616.5$ 1 852.91852.91\,852.91 852.9 $788.8$ 855 506855506855\,506855 506 0,1,2,4
C.2 Detailed Graphs with seeds
[…]
(a) 100 shots – $N{=}5$ qubits
(b) 100 shots – $N{=}10$ qubits
(c) 1000 shots – $N{=}5$ qubits
(d) 1000 shots – $N{=}10$ qubits

### v6

@@ C.6 Backward pass of hDQNN-TD3 (update phase)
On the same mini-batch we refine the surrogate by minimising the BCE loss of Eq. 5:
@@
Appendix D Algorithms
D.1 Constructive algorithm for the realised function Q
Algorithm 1 Evaluating the PQC layer $Q$ with $S$ shots, returning per-qubit empirical marginals
1:Input $\mathbf{q}_{i}$ (gate angles / routing), qubit count NN, depth $M$, shots $S$
2:Prepare $|\psi_{0}\rangle\leftarrow|0\rangle^{\otimes N}$
3:for $\ell=1$ to $M$ do
4:  for $i=1$ to NN do
5:   apply $R_{y}(\theta_{\ell,i}^{(y)})\,R_{z}(\theta_{\ell,i}^{(z)})$
6:  end for
7:  apply $\mathrm{CZ}$ on $(p_{\ell},k_{\ell})$ if $p_{\ell}\neq k_{\ell}$
8:end for
9:Measure all qubits in $Z$ for $S$ shots $\Rightarrow\{\mathbf{z}^{(s)}\}_{s=1}^{S}$
[…]
11:return $\hat{\mathbf{p}}\in[0,1]^{N}$ $\triangleright$ equivalently $\hat{\mathbf{m}}=1-2\hat{\mathbf{p}}$
D.2 Training loop pseudo-code for hDQNN-TD3
Algorithm 2 hDQNN-TD3: TD3 actor with a PQC and qtDNN surrogate
1:Environment $\mathcal{E}$, replay buffer $\mathcal{M}$, actor AA (PQC inside), twin critics $C_{1},C_{2}$, target networks $A^{\prime}$, $C_{1}^{\prime},C_{2}^{\prime}$, update periods $N_{\text{A}}$, surrogate epochs $N_{\text{qt}}$
2:for episode $1,2,\dots$ do
3:  collect experience: interact using $\mathbf{a}_{t}=A(\mathbf{s}_{t})+\mathcal{N}_{t}$, store $(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}},r_{t},\mathbf{s}_{t+1})$ in $\mathcal{M}$
4:  if time to update then
5:   Sample mini-batch $\mathcal{B}$ of size $N_{\text{b}}$ from $\mathcal{M}$
6:   (a) update qtDNN: train surrogate for $N_{\text{qt}}$ tiny-batches $\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}\in\mathcal{B}$
8:   if global step mod $N_{\text{A}}=0$ then
11:   end if
12:  end if
13:end for
Appendix E Additional graphs and tables
E.1 Model configuration
In this table, $\alpha$ is the learning rate and the Critic follows the structure from Lillicrap et al. [2019]
Table 2: hDQNN-TD3 Reinforcement Learning Model Configuration for Humanoid-v4 Problem
DNN Design Hyperparameters
Components Layers Activation NN,$M$ $d_{\text{c-link}}$ $\alpha$ $\gamma$ $\tau$ noise
PreDNN [376,256,230] LReLU 10 $3e^{-4}$ $5e^{-3}$
qtDNN [220,2 048,10] LReLU $3e^{-4}$
PQC 10,10 $1e^{-3}$
PostDNN [20,256,17] LReLU 10 $3e^{-4}$ $5e^{-3}$ $5e^{-2}$
Critic Block 1 [376,256] LReLU $3e^{-4}$ $5e^{-3}$
Critic Block 2 [273,256,1] LReLU $3e^{-4}$ 0.99 $5e^{-3}$
E.2 Extended Ablation
@@
4 0004\,000 FC 0 0 $423.0\pm 170.8$ $603.6$ $31.5$ 214 686214\,686 0,1,2,3,4
4 0004\,000 RBG 0 0 $333.6\pm 196.5$ $545.3$ $35.7$ 198 449198\,449 0,1,2,3,4
4 0004\,000 PQC $100$ 5 $1\,575.9\pm 1\,923.4$ 4 882.74\,882.7 $515.2$ 540 138540\,138 0,1,2,3,4
4 0004\,000 PQC 1 0001\,000 5 $476.5\pm 149.4$ $650.8$ $225.3$ 280 476280\,476 0,1,2,3,4
4 0004\,000 PQC $100$ 10 $667.3\pm 442.1$ 1 453.31\,453.3 $315.2$ 335 815335\,815 0,1,2,3,4
4 0004\,000 PQC 1 0001\,000 10 $556.0\pm 81.7$ $689.4$ $265.5$ 308 709308\,709 0,1,2,3,4
6 2006\,200 FC 0 0 $543.4\pm 243.3$ $909.4$ $72.2$ 423 010423\,010 0,1,2,3,4
6 2006\,200 RBG 0 0 $417.0\pm 131.4$ $526.2$ $75.5$ 366 904366\,904 0,1,2,3,4
6 2006\,200 PQC $100$ 5 $2\,014.7\pm 2\,008.6$ 4 342.44\,342.4 1 596.61\,596.6 1 387 7051\,387\,705 0,1,2,3,4
6 2006\,200 PQC 1 0001\,000 5 $544.8\pm 395.1$ 1 102.71\,102.7 $467.0$ 517 611517\,611 0,1,2,3,4
6 2006\,200 PQC $100$ 10 $1\,492.3\pm 2\,191.2$ 5 411.25\,411.2 $988.9$ 959 152959\,152 0,1,2,3,4
6 2006\,200 PQC 1 0001\,000 10 $657.0\pm 145.0$ $841.6$ $552.6$ 610 225610\,225 0,1,2,3,4
7 0007\,000 FC 0 0 $693.3\pm 453.1$ 1 473.01\,473.0 $98.5$ 515 399515\,399 0,1,2,3,4
7 0007\,000 RBG 0 0 $466.7\pm 106.0$ $547.2$ $92.8$ 438 296438\,296 0,1,2,3,4
7 0007\,000 PQC $100$ 5 2 $2\,139.6\pm 2\,180$.2 4 962.04\,962.0 1 935.91\,935.9 1 756 2661\,756\,266 0,1,2,3,4
7 0007\,000 PQC 1 0001\,000 5 $733.1\pm 740.4$ 1 974.31\,974.3 $612.5$ 642 867642\,867 0,1,2,3,4
7 0007\,000 PQC $100$ 10 $532.7\pm 105.6$ $690.6$ $678.9$ 666 056666\,056 0,1,2,4
7 0007\,000 PQC 1 0001\,000 10 $710.1\pm 238.6$ 1 040.61\,040.6 $675.5$ 744 615744\,615 0,1,2,3,4
7 2007\,200 FC 0 0 $617.5\pm 341.1$ 1 207.81\,207.8 $108.8$ 544 790544\,790 0,1,2,3,4
7 2007\,200 RBG 0 0 $414.3\pm 103.0$ $545.0$ $99.6$ 456 050456\,050 0,1,2,3,4
7 2007\,200 PQC $100$ 5 $1\,605.0\pm 2\,202.2$ 4 906.64\,906.6 1 791.91\,791.9 1 527 0371\,527\,037 0,1,2,4
7 2007\,200 PQC 1 0001\,000 5 $1\,078.5\pm 1\,517.3$ 3 759.23\,759.2 $673.0$ 692 138692\,138 0,1,2,3,4
7 2007\,200 PQC $100$ 10 $517.8\pm 95.0$ $659.2$ $712.2$ 694 104694\,104 0,1,2,4
7 2007\,200 PQC 1 0001\,000 10 $763.5\pm 335.7$ 1 277.41\,277.4 $711.9$ 782 646782\,646 0,1,2,3,4
7 7007\,700 FC 0 0 $1\,126.1\pm 1\,467.2$ 3 745.53\,745.5 $135.4$ 621 311621\,311 0,1,2,3,4
7 7007\,700 RBG 0 0 $432.4\pm 97.4$ $543.2$ $110.8$ 500 487500\,487 0,1,2,3,4
7 7007\,700 PQC $100$ 5 $1\,402.8\pm 1\,757.3$ 4 031.14\,031.1 1 943.21\,943.2 1 688 8881\,688\,888 0,1,2,4
7 7007\,700 PQC 1 0001\,000 5 $1\,121.4\pm 1\,558.9$ 3 868.63\,868.6 $839.3$ 824 076824\,076 0,1,2,3,4
7 7007\,700 PQC $100$ 10 $506.6\pm 97.0$ $624.5$ $793.6$ 765 043765\,043 0,1,2,4
7 7007\,700 PQC 1 0001\,000 10 $891.8\pm 547.1$ 1 852.91\,852.9 $805.0$ 882 668882\,668 0,1,2,3,4
@@
4 0004\,000 FC 0 0 $382.7\pm 167.5$ $603.6$ $27.4$ 196 401196\,401 0,1,2,4
4 0004\,000 RBG 0 0 $356.5\pm 219.1$ $545.3$ $39.8$ 210 584210\,584 0,1,2,4
4 0004\,000 PQC $100$ 5 $1\,555.4\pm 1\,923.4$ 4 882.74\,882.7 $564.8$ 574 890574\,890 0,1,2,4
4 0004\,000 PQC 1 0001\,000 5 $521.3\pm 128.0$ $650.8$ $237.5$ 302 738302\,738 0,1,2,4
4 0004\,000 PQC $100$ 10 $470.8\pm 57.1$ $534.5$ $291.2$ 302 948302\,948 0,1,2,4
4 0004\,000 PQC 1 0001\,000 10 $568.0\pm 89.2$ $689.4$ $258.4$ 303 057303\,057 0,1,2,4
6 2006\,200 RBG 0 0 $389.7\pm 134.3$ $521.2$ $80.4$ 372 226372\,226 0,1,2,4
6 2006\,200 FC 0 0 $514.3\pm 270.8$ $909.4$ $61.1$ 384 622384\,622 0,1,2,4
6 2006\,200 PQC $100$ 5 $1\,432.7\pm 1\,766.9$ 4 075.34\,075.3 1 489.31\,489.3 1 196 5031\,196\,503 0,1,2,4
6 2006\,200 PQC 1 0001\,000 5 $602.9\pm 430.8$ 1 102.71\,102.7 $498.4$ 567 036567\,036 0,1,2,4
6 2006\,200 PQC $100$ 10 $512.5\pm 52.6$ $590.5$ $551.2$ 557 415557\,415 0,1,2,4
6 2006\,200 PQC 1 0001\,000 10 $652.4\pm 167.0$ $841.6$ $544.0$ 593 357593\,357 0,1,2,4
7 0007\,000 FC 0 0 $691.6\pm 523.2$ 1 473.01\,473.0 $81.9$ 467 999467\,999 0,1,2,4
7 0007\,000 RBG 0 0 $446.5\pm 110.8$ $543.4$ $98.2$ 440 609440\,609 0,1,2,4
7 0007\,000 PQC $100$ 5 $1\,666.9\pm 2\,201.8$ 4 962.04\,962.0 1 730.91\,730.9 1 460 2281\,460\,228 0,1,2,4
7 0007\,000 PQC 1 0001\,000 5 $837.8\pm 811.1$ 1 974.31\,974.3 $664.2$ 711 615711\,615 0,1,2,4
7 0007\,000 PQC $100$ 10 $532.7\pm 105.6$ $690.6$ $678.9$ 666 056666\,056 0,1,2,4
7 0007\,000 PQC 1 0001\,000 10 $698.3\pm 273.8$ 1 040.61\,040.6 $665.6$ 724 890724\,890 0,1,2,4
7 2007\,200 FC 0 0 $620.8\pm 393.8$ 1 207.81\,207.8 $89.8$ 494 098494\,098 0,1,2,4
7 2007\,200 RBG 0 0 $426.8\pm 114.5$ $545.0$ $105.9$ 458 399458\,399 0,1,2,4
7 2007\,200 PQC $100$ 5 $1\,605.0\pm 2\,202.2$ 4 906.64\,906.6 1 791.91\,791.9 1 527 0371\,527\,037 0,1,2,4
7 2007\,200 PQC 1 0001\,000 5 1 $1\,269.5\pm 1\,681$.1 3 759.23\,759.2 $735.6$ 770 278770\,278 0,1,2,4
7 2007\,200 PQC $100$ 10 $517.8\pm 95.0$ $659.2$ $712.2$ 694 104694\,104 0,1,2,4
7 2007\,200 PQC 1 0001\,000 10 $768.9\pm 387.4$ 1 277.41\,277.4 $700.9$ 761 498761\,498 0,1,2,4
7 7007\,700 FC 0 0 1 $1\,253.1\pm 1\,662$.1 3 745.53\,745.5 $115.1$ 573 588573\,588 0,1,2,4
7 7007\,700 RBG 0 0 $431.2\pm 112.4$ $543.2$ $118.3$ 504 233504\,233 0,1,2,4
7 7007\,700 PQC $100$ 5 $1\,402.8\pm 1\,757.3$ 4 031.14\,031.1 1 943.21\,943.2 1 688 8881\,688\,888 0,1,2,4
7 7007\,700 PQC 1 0001\,000 5 $1\,326.7\pm 1\,720.3$ 3 868.63\,868.6 $933.4$ 927 753927\,753 0,1,2,4
7 7007\,700 PQC $100$ 10 $506.6\pm 97.0$ $624.5$ $793.6$ 765 043765\,043 0,1,2,4
7 7007\,700 PQC 1 0001\,000 10 $945.3\pm 616.5$ 1 852.91\,852.9 $788.8$ 855 506855\,506 0,1,2,4
E.3 Detailed Graphs with seeds
Figure 6 juxtaposes four learning curves obtained from an ablation study on the quantum layer’s hyper-parameters. Each subplot displays the $100$-episode moving-average return on Humanoid-v4 produced by the hDQNN-TD3 backbone while varying both the number of circuit shots, $\{100,\,1000\}$, and the number of active qubits, $\{5,\,10\}$. Every solid, colored trace corresponds to a single random seed; no additional smoothing or confidence band is applied, so the visible oscillations are the raw training dynamics. The comparison indicates that (i) increasing the qubit count from five to ten accelerates early learning only when the shot budget remains low, whereas (ii) allocating more shots generally stabilizes convergence but yields high returns chiefly in the smaller-qubit regime.
(a) 100 shots – N=5N{=}5 qubits
(b) 100 shots – N=10N{=}10 qubits
(c) 1000 shots – N=5N{=}5 qubits
(d) 1000 shots – N=10N{=}10 qubits